40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3882     Entry time: Mon Nov 8 18:30:33 2010
Author: Suresh 
Type: Update 
Category: IOO 
Subject: MC Trans Mon QPD gain increased by 50x 

Increased the transimpedance gain of the MC-Trans-Mon QPD ckt

The gain of this QPD was insufficient to see the light transmitted through the MC2.  The resulting voltage output was about  10 steps of the 16-bit ADC card.  As the input power, which is currently held at about 40mW may be increased to the vicinity of 2W (total output of the NPRO) we would have 500 ADC steps.  But the dynamic range of the ADC is 64k and  increasing the gain of this QPD ckt by a factor of 50 would enable us to utilise this dynamic range effectively.  However as we do not need a response faster than 10Hz from this ckt its response time has been limited by increasing the feedback capacitance value.

The ckt diagram for the QPD ckt is D980325-Rev-C1  .  The particular unit we are dealing with has the Serial No. 110.  The resistors R1, R2, R3, R4 are now 499 kOhm.  As per the guidelines in the ckt diagram, we increased the capacitance values C3,C4,C5,C6 to 2.2 nFThe current cut off frequency for the MC-Trans-Mon is 145 Hz (computed).

Initially, while reassembling the QPD unit, the IDC 16 connector to the ckt board was reversed by mistake and resulted in the OP497 chip over-heating.  Consequently one of the opamps on the chip was damaged and showed monotonously increasing ouput voltage.  Todd Etzel gave us a spare OP497 and I replaced the damaged chip with this new one. The chips are also available from  Newark Stock No. 19M8991 . The connector has been marked to indicate the correct orientation. The ckt was checked by temporarily connecting it in the place of the PRM Optical lever QPD.  It worked fine and has been put back in its place at the MC2 Transmission.  The QPD was wiped with a lens tissue+Methanol  to remove dust and finger prints from its surface.

It may need to be repositioned since the beam would have shifted under the MC realignment procedure.

 

 

 

ELOG V3.1.3-